1. Complete the table.

Substance Mass Moles

N0. of particle

(atoms or molecules)

Ar 5.0*10^-4
NO2 1.34*10^20
K 21.4 mg
C8H18 2.50 kg

2.Complete The table

Substance Mass Moles

Number OF particles

(atoms or molecules)

C6H12O6 19.0 g
Pb 8.24*10^21
CF4 24.5 kg
C 0.0406

3. Amylose is a "polysaccharide" that plants use to store energy. It is made of repeating subunits of C6H10O5.

a. A particular amylose molecule has 2113 of these subunits, what is its molecular formula? Express your answer as a chemical formula.

b. What is its molar mass? Express your answer to four significant figures and include the appropriate units.

c. What is the empirical formula for amylose? Express your answer as a chemical formula.

4. Imagine that all our balances displayed mass in units of ounces instead of grams.

Part A Would we still want to use the number , 6.022×1023, for Avogadro's number? Why or why not?

Yes, because the value of Avogadro's number is a universal constant.
Yes, because the number of particles in one mole does not depend on the mass of the particles.
No, because we would not be able to use carbon-12 to define the mole

No, because we would use the number of atoms in exactly 12 ounces of carbon-12 to define the mole.

Part B

What number would we want to use instead? (A ounce is equal to 28.350 g.)

Express your answer using four significant figures.

5. Hydrogen, a possible future fuel , can be obtained from other compounds such as ethanol. Ethanol can be made from the fermentation of crops such as corn.

Part A

How much hydrogen can be obtained from 5.0 kg of ethanol (C2H5OH)?

Express your answer using two significant figures.
Answer #1

1)
a)
Substance : Ar
molar mass of Ar = 40 g/mol

mass = molar mass*number of mol
= 40 g/mol * 5.0*10^-4 mol
= 0.02 g

number of particles = number of mol * Avogadro's number
= 5.0*10^-4 * 6.022*10^23
= 3.0*10^20 particles

b)
substance : NO2
1 NO2 has 3 particles

number of mol = number of particle / (3*Avogardro's numer)
moles = (1.34*10^20)/(3*6.022*10^23)
= 7.42*10^-5 mol

mass = number of mol * molar mass of NO2
= 7.42*10^-5 mol * 46 g/mol
= 3.41 g

c)
substance : K
moles = mass/molar mass of K
= 0.0214 g/39 g/mol
= 5.49*10^-4 mol

number of particle = 5.49*10^-4 * 6.022*10^23
= 3.30*10^20

d)
substance : C8H18
moles = mass/molar mass of C8H18
= 2.50*10^3 g/114 g/mol
= 21.9 mol

number of particle = 21.9* 26*6.022*10^23
= 3.43*10^26

I am allowed to answer only 1 question at a time